# Model-independent predictions for decays of double-heavy hadrons into pairs of heavy hadrons

### v2

@@ V.3 Relative Partial Decay Rates
The transition amplitudes in Eq. (63) depend on the spatial overlap of the quark-model wavefunctions only through the matrix elements $A$. The sums in Eq. (63) run over the two orbital-angular-momentum quantum numbers $L^{\prime}$ and $L^{\prime\prime}$. If the sums reduce to a single term, there is only one dominant spatial matrix element $A$ so it factors out of transition amplitudes of quarkonium states with the same quantum numbers $n$, $L_{Q}$ and all possible values of $S_{Q}$, $J$ into pairs of mesons with the same quantum numbers $n_{1}$, $n_{2}$, $L_{1}$, $L_{2}$, $L_{Q}^{\prime}$ and all possible values of $S_{1}$, $S_{2}$, $J_{1}$, $J_{2}$. In this case, nonflip triplet decay models can be used to predict the corresponding relative partial decay rates under the conditions outlined in Section III.3.
@@
Note that the B-O approximation is generally more predictive than nonflip triplet decay models because it incorporates cylindrical and heavy-quark spin symmetries. That is exemplified by the decays of $P$-wave quarkonia into one $S$-wave meson and one $P$-wave meson in a relative $P$-wave. In nonflip triplet decay models, the relative partial decay rates are not completely determined by symmetries since these decays do not fall into either case a) or b). The B-O symmetries, on the other hand, fully constrain the relative partial decay rates for $P$-wave quarkonia into $S+P^{-}$ ($P$-wave) and $S+P^{+}$ ($P$-wave) to the values listed in Table 2.
While the B-O approximation reproduces relative partial decay rates from quark-pair-creation models of quarkonium decays, we find no support for such models of quarkonium hybrid decays. For quarkonium hybrids, the B-O angular-momentum coefficients and relative partial decay rates differ from those of nonflip-triplet decay models even in simple cases. This disagreement is in addition to the already mentioned violation of the quark-pair-creation-model selection rule that prohibits decays of quarkonium hybrids into pairs of mesons with identical spatial wavefunctions.
@@ VI Overview
In this work, we have derived model-independent results for the decays of double-heavy hadrons into pairs of heavy hadrons using the diabatic B-O approximation for QCD. Our general expressions for the coupling potentials between a double-heavy hadron and heavy-hadron pairs in Eq. (41) are sums of products of B-O transition amplitudes and angular-momentum coefficients. The B-O transition amplitudes $g_{\lambda,\eta}$ can be calculated using lattice QCD with two static color sources. The angular-momentum coefficients in Eq. (43) are determined by products of Clebsch-Gordan coefficients, Wigner 6-$j$ symbols, and Wigner 9-$j$ symbols.
We have used these expressions to derive selection rules for the decays of double-heavy hadrons into pairs of heavy hadrons. The selection rules are given in Section III.2. We also obtained analytic approximations for relative partial decay rates between double-heavy hadrons in the same B-O multiplet and pairs of heavy hadrons in specific heavy-quark-spin-symmetry doublets. The relative partial decay rates are given in Eq. (54). Our results reproduce and extend previous ones obtained for quarkonium decays using heavy-quark spin symmetry alone in Refs. [24, 25, 26]. Model-independent results have been obtained previously for exotic heavy mesons under the assumption that they are heavy-meson molecules [39, 40, 41, 42, 43]. Our methods can be used to extend those results.
@@
It is important to emphasize that the results of this paper do not apply exclusively to the decays of quarkonia and quarkonium hybrids into pairs of heavy mesons. They can be equally well applied to their decays into a pair of heavy baryons, the decays of a double-heavy baryon into a heavy meson and a heavy baryon, and the decays of a double-heavy tetraquark into a pair of heavy mesons or a pair of heavy baryons. All it takes is to identify the B-O quantum numbers of the decaying double-heavy hadron and specify the light-QCD quantum numbers of the two heavy hadrons in the final state. One can derive the corresponding selection rules by applying our general expressions in Section III.2. If the decays are dominated by a single B-O transition amplitude, the ratio of partial decay rates is given by our general analytic expression in Eq. (54).
The expressions for the coupling potentials obtained by inserting Eq. (41) into Eq. (43) can be combined with the techniques pioneered in Ref. [21] to include spin splittings of heavy hadrons in the diabatic B-O framework for double-heavy hadrons coupled with heavy-hadron pairs. This is extremely important for exotic double-heavy hadrons, since being able to resolve heavy-hadron spin splittings grants access to the near-threshold dynamics that characterizes many exotic states. Using the techniques of Ref. [27] together with those of this paper, calculating the spectrum of both conventional and exotic double-heavy hadrons boils down to calculating just a handful of B-O potentials and transition amplitudes using lattice QCD with two static color sources. We strongly encourage theoretical efforts in this direction.

### v3

@@ V.3 Relative Partial Decay Rates
The transition amplitudes in Eq. (50) depend on the spatial overlap of the quark-model wavefunctions only through the matrix elements $A$. The sums in Eq. (50) run over the two orbital-angular-momentum quantum numbers $L^{\prime}$ and $L^{\prime\prime}$. If the sums reduce to a single term, there is only one dominant spatial matrix element $A$ so it factors out of transition amplitudes of quarkonium states with the same quantum numbers $n$, $L_{Q}$, and all possible values of $S_{Q}$, $J$ into pairs of mesons with the same quantum numbers $n_{1}$, $n_{2}$, $L_{1}$, $L_{2}$, $L_{Q}^{\prime}$, and all possible values of $S_{1}$, $S_{2}$, $J_{1}$, $J_{2}$. In this case, nonflip triplet decay models can be used to predict the corresponding relative partial decay rates under the conditions outlined in Section III.3.
@@
Note that the B-O approximation is generally more predictive than nonflip triplet decay models because it incorporates cylindrical symmetries. That is exemplified by the decays of $P$-wave quarkonia into one $S$-wave meson and one $P$-wave meson in a relative $P$-wave. In nonflip triplet decay models, the relative partial decay rates are not completely determined by symmetries since these decays do not fall into either case a) or b). The B-O symmetries, on the other hand, fully constrain the relative partial decay rates for $P$-wave quarkonia into $S+P^{-}$ ($P$-wave) and $S+P^{+}$ ($P$-wave) to the values listed in Table 2.
While the B-O approximation reproduces relative partial decay rates from nonflip-triplet models of quarkonium decays, we find no support for such models of quarkonium hybrid decays. For quarkonium hybrids, the B-O angular-momentum coefficients and relative partial decay rates differ from those of nonflip-triplet decay models even in simple cases. This disagreement is in addition to the already mentioned violation of the quark-pair-creation-model selection rule that prohibits decays of quarkonium hybrids into pairs of mesons with identical spatial wavefunctions.
@@ VI Overview
In this work, we have derived model-independent results for the decays of hidden-heavy hadrons into pairs of heavy hadrons using the diabatic B-O approximation for QCD. Our general expressions for the coupling potentials between a hidden-heavy hadron and heavy-hadron pairs in Eq. (30) are sums of products of B-O transition amplitudes and angular-momentum coefficients. The B-O transition amplitudes $g_{\lambda,\eta}$ can be calculated using lattice QCD with static $\bm{3}$ and $\bm{3}^{\ast}$ color sources. The angular-momentum coefficients in Eq. (32) are determined by products of Clebsch-Gordan coefficients, Wigner 6-$j$ symbols, and Wigner 9-$j$ symbols.
We have used these expressions to derive selection rules for the decays of hidden-heavy hadrons into pairs of heavy hadrons. The selection rules are given in Section III.2. We also obtained analytic approximations for relative partial decay rates between hidden-heavy hadrons in the same B-O multiplet and pairs of heavy hadrons in specific heavy-quark-spin-symmetry doublets. The relative partial decay rates are given in Eq. (41). Our results reproduce and extend previous ones obtained for quarkonium decays using heavy-quark spin symmetry alone in Refs. [24, 25, 26]. Model-independent results have been obtained previously for exotic hidden-heavy mesons under the assumption that they are heavy-meson molecules [40, 41, 42, 43, 44]. Our methods can be used to extend those results.
@@
It is important to emphasize that the results of this paper do not apply exclusively to the decays of quarkonia and quarkonium hybrids into pairs of heavy mesons. They can be equally well applied to their decays into a pair of heavy baryons. They can also be applied to the decays of other exotic hidden-heavy hadrons, including hidden-heavy tetraquark mesons and hidden-heavy pentaquark baryons. All it takes is to identify the B-O quantum numbers of the decaying hidden-heavy hadron and to specify the light-QCD quantum numbers of the two heavy hadrons in the final state. One can derive the corresponding selection rules by applying our general expressions in Section III.2. If the decays are dominated by a single B-O transition amplitude, the ratio of partial decay rates is given by our general analytic expression in Eq. (41).
Similar methods can be applied to the decays of double-heavy hadrons into pairs of heavy hadrons. A double-heavy hadron contains two heavy quarks with the same flavor. The B-O transition amplitude for its decays into pairs of heavy hadrons can be calculated using lattice QCD with two static $\bm{3}$ color sources. The selection rules and relative partial decay rates for decays of double-heavy hadrons are more intricate than those for decays of hidden-heavy hadrons, because light flavor symmetries play a more significant role [46].
The expressions for the coupling potentials obtained by inserting Eq. (30) into Eq. (32) can be combined with the techniques pioneered in Ref. [21] to include spin splittings of heavy hadrons in the diabatic B-O framework for hidden-heavy hadrons coupled with heavy-hadron pairs. That is extremely important for exotic hidden-heavy hadrons, since being able to resolve heavy-hadron spin splittings grants access to the near-threshold dynamics that characterizes many exotic states. Using the techniques of Ref. [21] together with those of this paper, calculating the spectrum of both conventional and exotic hidden-heavy hadrons boils down to calculating just a handful of B-O potentials and transition amplitudes using lattice QCD with static $\bm{3}$ and $\bm{3}^{\ast}$ color sources. We strongly encourage theoretical efforts in this direction.